the length of rectangle is 2cm more then twice its breadth . if the perimeter of rectangle 28cm.fins its length and breadth
Answer by fractalier(6550)

Call length and breadth, L and B.
Thus L = 2B + 2
Also perimeter P = 2L + 2B = 28
Substitute the first equation into the second and get
2(2B + 2) + 2B = 28
4B + 4 + 2B = 28
6B + 4 = 28
6B = 24
B = 4
L = 10